The amplitude of a lightly damped oscillator decreases by 3.0% during each cycle.After how many oscillations(approximately) will the amplitude drop to its half its initial value


1
Expert's answer
2021-04-09T06:34:40-0400

"\frac A2=A(1-\frac{k}{100})^n,"

"\frac 12=(1-\frac{k}{100})^n,"

"n=\frac{1}{log_2\frac{100}{100-k}}=23."